OpenStudy (anonymous):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

hero (hero):

Hint: 2(2y - 1) + y = 3

OpenStudy (anonymous):

I don't know lol

OpenStudy (jennychan12):

plug in x for x in the first equation. 2(2y-1) +y = 3 now that you have everything in terms of y solve for y then plug in y for x = 2y-1

OpenStudy (jennychan12):

4y - 2 +y = 3 5y = 5 y =1 x = 2-1 x=1